write the equation of a line that has a slope of 5/2 and passes through (2,3).
Answer by nerdybill(7384)

write the equation of a line that has a slope of 5/2 and passes through (2,3).
.
Plug given info into the "point-slope" form:
y - y1 = m(x - x1)
y - 3 = (5/2)(x - 2)
y - 3 = (5/2)x - 5
y = (5/2)x - 2 (this is what they're looking for)